# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 5.3 The maximum of the maximum
Before we compute its optimal solution for a specific choice of $\mu_{T}$, we will have a look at the corresponding continuous-time solution. Fix a probability space $(\Omega,\mathcal{F},\mathbb{Q})$ and let $S^{\text{AY}}$ denote the martingale on $[0,T]$ whose maximum has a first-order stochastic dominance over the maximum of any other continuous martingale with terminal marginal $\mu_{T}$. It exists, and is related to the martingale constructed by Azéma and Yor [3] as a solution of the Skorokhod embedding problem with terminal marginal $\mu_{T}$ [10, Lemma 2.2]; we will refer to this martingale as the Azéma–Yor martingale. It follows that for the continuous-time case, an optimal solution of the above MOT problem is given by the Azéma–Yor martingale $S^{\text{AY}}$ [38]. A solution of the problem in continuous-time is thus known. It follows from [11, Lemma 2.4] that said distribution can be obtained as
@@
(b) Corresponding function $f_{B}(y)$.
Figure 3: The given terminal marginal used in Example 5.4 is displayed in Figure 3(a). It is centered in $0.5$ and its support is contained within $[0,1]$. The function $f_{B}$ is shown in Figure 3(b) for a few different values of $B$.
Let the terminal marginal $\mu_{T}$ be as in Figure 3(a) with $s_{0}=0.5$, see Figure 3(b) for examples of the function $f_{B}$ corresponding to $\mu_{T}$ for a few different values of $B$ — note that its minimum exists for each $B$. The probability $\mathbb{Q}(X_{T}^{\text{AY}}\geq B)$ is obtained by minimising $f_{B}$ over $[0,B]$. By doing this for all $B\in(0.5,1]$ we are able to recover the distribution of $X_{T}^{\text{AY}}$. It is displayed using a bold red line in Figure 4(a). The corresponding problem value, the robust price, is $\mathbb{E}_{\mathbb{Q}}\left[X_{T}^{\text{AY}}\right]=0.6133$.
We now solve the corresponding upper bound MOT problem for different values of $T$, ranging from $2$ to 29, by application of our framework with $\varepsilon=0.01$. The corresponding law of the maximum of the price process is displayed in Figure 4(a). The corresponding robust prices, $\mathbb{E}_{\mathbb{Q}}[X_{T}]$, are shown in Figure 4(b) as a function of the number of marginals used. We note that the computed distribution approaches the known continuous-time solution and that the cumulative distribution function decreases as the number of time steps used increases. Similarly the computed problem values, the robust prices, approaches the problem value of the corresponding continuous-time problem.
(a) Computed cumulative distribution functions of the maximum of the maximum for an increasing number of marginals.
(b) The computed robust price as a function of the number of marginals used.
Figure 4: The distribution of the maximum of the maximum from Example 5.4, solved approximately for an increasing number of marginals $T+1$, is shown in Figure 4(a). The corresponding distribution for the continuous-time case is included for reference (bold red). We note that the discrete-time computed solutions approach the continuous-time solution as the number of time steps used increases, and that the cumulative distribution function decreases. The corresponding problem values are shown as a function of the number of marginals $T+1$ in Figure 4(b); the problem value for the continuous-time case is included for comparison (bold red).
@@ 5.4 The robust price of a digital option
$\mathbb{E}_{\mathbb{Q}}\left[\chi_{[B,\infty)}\left(\underset{0\leq t\leq T}{\text{max }}S_{t}^{\text{AY}}\right)\right]\geq\mathbb{E}_{\mathbb{Q}}\left[\chi_{[B,\infty)}\left(\underset{0\leq t\leq T}{\text{max }}S_{t}\right)\right],\quad B>s_{0}.$
For a specific digital option though — that is, for $B$ fixed — one can find an (optimal) coupling with the same objective value by using just one intermediate time step [24, pp.416–419]. In the below examples, we solve problems with $T=2$ computationally; we will recover the coupling constructed in [24] as well as the distribution of the maximum of the Azéma–Yor martingale.
@@
Let $s_{0}=0.5$ and $\mu_{T}=0.5(\delta_{0}+\delta_{1})$, then it follows from equation 39 that the the robust price of a digital option with barrier $B\in(0.5,1]$ for the continuous-time problem is $(2B)^{-1}$. On the other hand, [24, Theorem 7.27] yields the same value when taking only discrete time models with $T\geq 2$ into account. Repeating the construction from the proof of the theorem allows for constructing a martingale coupling that attains this upper bound. For this example with $T=2$, it corresponds to take at $t=1$ any of the two values 0 or $B$ with probability $1-(2B)^{-1}$ and $(2B)^{-1}$, respectively, before moving to the terminal marginal. We note that the three marginals fully specifies a martingale coupling.
We approximately solve the corresponding tri-marginal MOT problem by application of our framework for $B=0.75$, $\varepsilon=0.02$ and $n_{S}=100$. The corresponding intermediate marginal distribution ($t=1$) of the price process is $0.27\delta_{0}+0.05\delta_{0.01}+0.01\delta_{0.02}+0.67\delta_{0.75}$; this can be compared with $0.33\delta_{0}+0.67\delta_{0.75}$, which is given by the theoretical construction. The computed distribution is subject to some smoothing around zero, which is due to regularisation.
@@
Let $T=2$ and let the initial and terminal marginals be as in Example 5.4. We solve the corresponding MOT problem computationally for a set of barriers ranging from 0.5 to 0.92; a regularisation parameter of $\varepsilon=0.01$ was used. By doing so we obtain the maximum of the value $\mathbb{E}_{\mathbb{Q}}[\chi_{[B,\infty)}(\max_{t\in[T]}S_{t})]=\mathbb{Q}(\max_{t\in[T]}S_{t}\geq B)$ for each such barrier $B$; the result is shown as a function of the barrier in Figure 5, along with the cumulative distribution function of the maximum of $S^{\text{AY}}$ (cf. Example 5.4). We see that by successively solving for the upper bounds on the probability that digital options with barriers $B\in(0.5,1]$ are not activated, we can approximately reproduce the cumulative distribution function of the maximum of the Azéma–Yor martingale.
Figure 5: The computed upper bound on the probability that the digital option with barrier $B$ is not activated as a function of the barrier level $B$, as given in Example 5.6. The cumulative distribution function of the maximum of the corresponding Azéma–Yor martingale is displayed in bold red for reference. The grey vertical lines illustrate the support of the price process contained within $[0.5,0.92]$. Note that the resolution of the computed approximations follow this grid.
@@ 5.5 Robust pricing of an Asian option
We conclude by considering an example where the optimal solution is not known, namely, the robust pricing of an Asian option subject to several marginal constraints. Asian options are a class of financial derivatives that are of great practical financial interest; they are characterised by their payoff being of the form
$\phi(S_{0},\dots,S_{T})=f\bigg{(}\frac{1}{T+1}\sum_{t\in[T]}S_{t}\bigg{)}$
for some function $f:\mathbb{R}\rightarrow\mathbb{R}$. The dependence on the arithmetic mean makes the corresponding MOT problem a difficult one to solve analytically — this was discussed in [57], who provided the optimal solutions for a problem with two or three given marginals and $f$ convex. As far as we know, it is still unclear what the optimal solution would look like for a MOT problem subject to more than four marginal constraints. By letting the stochastic process $X$ be as given in Example 2.2, we can however solve these problems computationally. We will now do so for an example where we successively increase the number of marginal constraints. In order to provide visually clear results, all given marginals will be rather simple.
@@
Let the payoff function $\phi$ be as given above with $f(x)=|x-30|$. Then the payoff corresponds to the payoff of an Asian straddle with strike 30. Let $T=11$. Consider a trivial initial marginal $\mu_{0}=\delta_{30}$ and a uniformly distributed terminal marginal $\mu_{11}$ whose support is contained within $n_{S}=41$ atoms, evenly distributed over the interval $[25,35]$. In order not to run into problems with rounding errors, we will here use an increasing grid for representing the support of the process $X$; it is straightforward to generalise the results from Section 4 to this setting.
The optimal solution of the corresponding lower bound MOT problem, subject to marginal constraints on the initial and the terminal marginals, was computed using $\varepsilon=0.006$. The marginals of the computed optimal coupling are displayed in Figure 6(a), where the marginals that are subject to a constraint are marked in green. We note that it corresponds to a late transport, as given by [57]. The smoothing is due to regularisation. A third marginal constraint was then added on the fourth marginal, by requiring that $\mu_{4}=\frac{1}{2}(\delta_{29}+\delta_{31})$. The problem was then solved again; the marginals of the computed optimal coupling are shown in Figure 6(b). We note that the transport from the initial marginal $\mu_{0}$ to the given intermediate marginal $\mu_{4}$ corresponds to a late transport, as conjectured by [57]. The problem was then solved a third time with a fourth marginal constraint imposed on the eighth marginal, with $\mu_{8}=\frac{1}{4}(\delta_{28}+2\delta_{30}+\delta_{32})$. The marginals of the optimal coupling are shown in Figure 6(c).
(a) $\mathcal{T}=\{0,11\}$
(b) $\mathcal{T}=\{0,4,11\}$
(c) $\mathcal{T}=\{0,4,8,11\}$

### v2

@@ 5.3 The maximum of the maximum
Before we compute its optimal solution for a specific choice of $\mu_{T}$, we will have a look at the corresponding continuous-time solution. Fix a probability space $(\Omega,\mathcal{F},\mathbb{Q})$ and let $S^{\text{AY}}$ denote the martingale on $[0,T]$ whose maximum has a first-order stochastic dominance over the maximum of any other continuous martingale with terminal marginal $\mu_{T}$. It exists, and is related to the martingale constructed by Azéma and Yor [3] as a solution of the Skorokhod embedding problem (SEP) with terminal marginal $\mu_{T}$ [12, Lemma 2.2]; we will refer to this martingale as the Azéma–Yor martingale. It follows that for the continuous-time case, an optimal solution of the above MOT problem is given by the Azéma–Yor martingale $S^{\text{AY}}$ [43]. A solution of the problem in continuous-time is thus known. It follows from [13, Lemma 2.4] that said distribution can be obtained as
@@
Let the terminal marginal $\mu_{T}$ be as in Figure 3(a) with $s_{0}=0.5$ and $n_{T}^{\mathcal{S}}=67$. The probability $\mathbb{Q}(X_{T}^{\text{AY}}\geq B)$ is obtained by minimising $f_{B}$ over $[0,B]$ — note that its minimum exists for each $B$. By doing this for all $B\in(0.5,1]$ we are able to recover the distribution of $X_{T}^{\text{AY}}$. It is displayed using a bold red line in Figure 3(b). The corresponding problem value, the robust price, is $\mathbb{E}_{\mathbb{Q}}[X_{T}^{\text{AY}}]=0.6133$.
We now solve the corresponding upper bound MOT problem for different values of $T$, ranging from $2$ to 29, by application of our framework with $\varepsilon=0.01$. The corresponding law of the maximum of the price process is displayed in Figure 3(b). The corresponding robust prices, $\mathbb{E}_{\mathbb{Q}}[X_{T}]$, are shown in Figure 3(c) as a function of the number of marginals used. We note that the computed distribution approaches the known continuous-time solution and that the cumulative distribution function decreases as the number of time steps used increases. Similarly the computed problem values, the robust prices, approaches the problem value of the corresponding continuous-time problem.
For the problem with $T=29$ we computed the residuals of all constraints, prior to updating the corresponding dual variable. The maximum element of the residuals is shown in Figure 3(d) as a function of the Sinkhorn iteration count, for both marginal constraints and for some of the martingale constraints, along with the threshold level used in Newton’s method. Note that a high level of accuracy was used in this example — the tolerated marginal and martingale residuals were set to $10^{-6}$ and $10^{-8}$, respectively. This is why the number of iterations in Figure 3(d) is quite large.
@@
(b) Computed cumulative distribution functions of the maximum of the maximum for an increasing number of marginals.
(c) The computed robust price as a function of the number of marginals used.
(d) Computed residuals as a function of the Sinkhorn iteration count when $T=29$.
Figure 3: The given terminal marginal used in Example 5.4 is displayed in Figure 3(a). It is centered in $0.5$ and its support is contained within $[0,1]$. The corresponding distribution of the maximum of the maximum, solved approximately for an increasing number of marginals $T+1$, is shown in Figure 3(b). The corresponding distribution for the continuous-time case is included for reference (bold red). We note that the discrete-time computed solutions approach the continuous-time solution as the number of time steps used increases, and that the cumulative distribution function decreases. The corresponding problem values are shown as a function of the number of marginals $T+1$ in Figure 3(c); the problem value for the continuous-time case is included for comparison (bold red). The maximum element of the computed residuals, obtained prior to updating the corresponding dual variable, for the marginal constraints (dashed line) and for some of the martingale constraints (bold line) are displayed in a linlog scale in Figure 3(d), along with the threshold level used in Newton’s method (black dashed line). Note the high level of accuracy used in the computations.
Remark 5.5 (Other applications to computation of optimal Skorokhod embeddings).
In the above we saw that the Azéma-Yor martingale maximises the law of the maximum among all continuous martingales that respects the terminal law $\mu_{T}$, and that our framework can be used to approximately obtain this law by computing an upper bound MOT in discrete-time for $T$ large. There is another martingale, constructed from the Perkins solution of the same SEP, that minimises the law of the maximum among all continuous martingales that has terminal law $\mu_{T}$. We could analogously approximately compute this minimal law by solving the corresponding lower bound MOT problem.
If we instead let the auxiliary process $X$ be the realised variance of the price process $S$ (cf. Example 2.5) we can in a similar manner for $T$ large approximate the martingale that is induced by yet another solution of the SEP with terminal marginal $\mu_{T}$, namely, the Root solution. Indeed, it follows from results by Dupire and by Carr and Lee that it solves the corresponding continuous-time lower bound MOT problem with $\phi(S_{0},\dots,S_{T})=f(X_{T})$ for $f:\mathbb{R}\rightarrow\mathbb{R}$ convex. See [42] for further details.
@@ 5.4 The robust price of a digital option
$\mathbb{E}_{\mathbb{Q}}\big{[}\chi_{[B,\infty)}\big{(}\max_{t\in[0,T]}S_{t}^{\text{AY}}\big{)}\big{]}\geq\mathbb{E}_{\mathbb{Q}}\big{[}\chi_{[B,\infty)}\big{(}\max_{t\in[0,T]}S_{t}\big{)}\big{]},\quad B>s_{0}.$
For a specific digital option though — that is, for $B$ fixed — one can find an (optimal) coupling with the same objective value by using just one intermediate time step [27, pp. 416–419]. In the below examples, we solve problems with $T=2$ computationally; we will recover the coupling constructed in [27] as well as the distribution of the maximum of the Azéma–Yor martingale.
@@
Let $s_{0}=0.5$ and $\mu_{T}=0.5(\delta_{0}+\delta_{1})$, then it follows from equation 35 that the the robust price of a digital option with barrier $B\in(0.5,1]$ for the continuous-time problem is $(2B)^{-1}$. On the other hand, [27, Theorem 7.27] yields the same value when taking only discrete time models with $T\geq 2$ into account. Repeating the construction from the proof of the theorem allows for constructing a martingale coupling that attains this upper bound. For this example with $T=2$, it corresponds to take at $t=1$ any of the two values 0 or $B$ with probability $1-(2B)^{-1}$ and $(2B)^{-1}$, respectively, before moving to the terminal marginal. We note that the three marginals fully specifies a martingale coupling.
We approximately solve the corresponding tri-marginal MOT problem by application of our framework for $B=0.75$, $\varepsilon=0.02$ and $n_{1}^{\mathcal{S}}=n_{2}^{\mathcal{S}}=100$. The corresponding intermediate marginal distribution ($t=1$) of the price process is $0.27\delta_{0}+0.05\delta_{0.01}+0.01\delta_{0.02}+0.67\delta_{0.75}$; this can be compared with $0.33\delta_{0}+0.67\delta_{0.75}$, which is given by the theoretical construction. The computed distribution is subject to some smoothing around zero, which is due to regularisation.
@@
Let $T=2$ and let the initial and terminal marginals be as in Example 5.4. We solve the corresponding MOT problem computationally for a set of barriers ranging from 0.5 to 0.92; a regularisation parameter of $\varepsilon=0.01$ was used and a number of $n_{1}^{\mathcal{S}}=n_{2}^{\mathcal{S}}=67$ points of support. By doing so we obtain the maximum of the value $\mathbb{E}_{\mathbb{Q}}[\chi_{[B,\infty)}(\max_{t\in[T]}S_{t})]=\mathbb{Q}(\max_{t\in[T]}S_{t}\geq B)$ for each such barrier $B$; the result is shown as a function of the barrier in Figure 4, along with the cumulative distribution function of the maximum of $S^{\text{AY}}$ (cf. Example 5.4). We see that by successively solving for the upper bounds on the probability that digital options with barriers $B\in(0.5,1]$ are not activated, we can approximately reproduce the cumulative distribution function of the maximum of the Azéma–Yor martingale.
Figure 4: The computed upper bound on the probability that the digital option with barrier $B$ is not activated as a function of the barrier level $B$, as given in Example 5.7. The cumulative distribution function of the maximum of the corresponding Azéma–Yor martingale is displayed in bold red for reference. The grey vertical lines illustrate the support of the price process contained within $[0.5,0.92]$. Note that the resolution of the computed approximations follows this grid.
@@ 5.5 Robust pricing of an Asian option
We conclude by considering an example where the optimal solution is not known, namely, the robust pricing of an Asian option subject to several marginal constraints. Asian options are a class of financial derivatives that are of great practical financial interest; they are characterised by their payoff being of the form $\phi(S_{0},\dots,S_{T})=f((T+1)^{-1}\sum_{t\in[T]}S_{t})$ for some function $f:\mathbb{R}\rightarrow\mathbb{R}$. The dependence on the arithmetic mean makes the corresponding MOT problem a difficult one to solve analytically — this was discussed in [63], who provided the optimal solutions for a problem with two or three given marginals and $f$ convex. As far as we know, it is still unclear what the optimal solution would look like for an MOT problem subject to more than four marginal constraints. By letting the stochastic process $X$ be as given in Example 2.3, we can however solve these problems computationally. We will now do so for an example where we successively increase the number of marginal constraints. In order to provide visually clear results, all given marginals will be rather simple.
@@
Let the payoff function $\phi$ be as given above with $f(x)=|x-30|$. Then the payoff corresponds to the payoff of an Asian straddle with strike 30. Let $T=11$. Consider a trivial initial marginal $\mu_{0}=\delta_{30}$ and a uniformly distributed terminal marginal $\mu_{11}$ whose support is contained within $n_{11}^{\mathcal{S}}=41$ atoms, evenly distributed over the interval $[25,35]$. In order not to run into problems with rounding errors, we will here use an increasing grid for representing the support of the process $X$; we use a number of points of support ranging from $n_{0}^{\mathcal{X}}=1$ to $n_{11}^{\mathcal{X}}=481$ while $n_{t}^{\mathcal{S}}=41$ for $t\geq 1$.
The optimal solution of the corresponding lower bound MOT problem, subject to marginal constraints on the initial and the terminal marginals, was computed using $\varepsilon=6\cdot 10^{-3}$. The marginals of the computed optimal coupling are displayed in Figure 5(a), where the marginals that are subject to a constraint are marked in green. We note that it corresponds to a late transport, as given by [63]. The smoothing is due to regularisation. A third marginal constraint was then added on the fourth marginal, by requiring that $\mu_{4}=\frac{1}{2}(\delta_{29}+\delta_{31})$. The problem was then solved again; the marginals of the computed optimal coupling are shown in Figure 5(b). We note that the transport from the initial marginal $\mu_{0}$ to the given intermediate marginal $\mu_{4}$ corresponds to a late transport, as conjectured by [63]. The problem was then solved a third time with a fourth marginal constraint imposed on the eighth marginal, with $\mu_{8}=\frac{1}{4}(\delta_{28}+2\delta_{30}+\delta_{32})$. The marginals of the optimal coupling are shown in Figure 5(c).